OpenStudy (anonymous):

What is 2x + 4 if x = 14

hartnn (hartnn):

just put x=14 there 2(14) +4 = ?

OpenStudy (anonymous):

6x

hartnn (hartnn):

no! when u put x=14 u get 2(14)+4 what is 2(14) = ?

OpenStudy (anonymous):

6x?

hartnn (hartnn):

IT IS NOT 6x

OpenStudy (anonymous):

What I found it as 6x

hartnn (hartnn):

how 6 ?

OpenStudy (anonymous):

Because 14*2 == 6

hartnn (hartnn):

14*2=28

OpenStudy (anonymous):

What Liar IT is 6x

OpenStudy (anonymous):

.. your math is flawed. also, where did you get an x from? when you replace the x, there is no more x. 14 is in place of the x

OpenStudy (anonymous):

@DrTacoPhD I know more than you I do not need you help it is 6x

OpenStudy (anonymous):

if you know more, why bother asking the question?